OpenStudy (anonymous):

the problem 1A-4 c) don't understand why profesor cancel 2 2a/a^2-x^2 - 2x/a^2-x^2 and he end with a/a^2-x^2 - x/a^2-x^2 why he cancel 2 at the numerator please if somebody wi'll help me

OpenStudy (anonymous):

If you look at it,if we sum up the two equations [1/x+a+1/-x+a\] and \[1/x+a-1/-x+a\] on the left-hand side you get [2/x+a\] (1/-x+a and -1/-x+a cancel out) and on the right hand side there is 2a/a^{2}-x^{2} -2x/a^{2}-x^{2}=2*(a-x)/(a^{2}-x^{2}). So there is a 2 on each side,and they cancel each other leaving just the final equation

OpenStudy (anonymous):

can you explain in it better? is this the equation? \[\frac{2a }{ a ^{2}-x ^{2}}-\frac{ 2x }{ a ^{2}-x ^{2} }=\frac{ a }{ a ^{2}-x ^{2}}-\frac{ x }{ a ^{2}-x ^{2}}\]